An average emf of 20 V is induced in an inductor when the current in it is changed from 2.5 A in one direction to the same value in the opposite direction in 0.1 s. Find the self-inductance of the inductor.

Grade:10

1 Answers

Navjyot Kalra
9 years ago
Sol. V = 20 V dI = I base 2 – I base 1 = 2.5 – (–2.5) = 5A dt = 0.1 s V = L dl/dt ⇒ 20 = L(5/0.1) ⇒ 20 = L × 50 ⇒ L = 20 / 50 = 4/10 = 0.4 Henry.